OpenStudy (anonymous):

simplify complex fractions:

OpenStudy (anonymous):

\[x ^{-2}-9y ^{-2} \over 5y-15x\]

OpenStudy (anonymous):

\[\frac{\frac{1}{x^2}-\frac{9}{y^2}}{5y-12x}\] is a start. then perhaps multiply top and bottom by \(x^2y^2\) to clear the fraction

OpenStudy (anonymous):

thats 15 not12

OpenStudy (anonymous):

you should get \[\frac{y^2-9x^2}{x^2y^2(5y-15x)}\] then factor and cancel a bit

OpenStudy (anonymous):

\[\frac{(y-3x)(y+3x)}{5x^2y^2(y-3x)}\] one step left

OpenStudy (anonymous):

um cancel out y-3x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw